35. A basketball is inflated to a pressure of 1.50 atm in a 20.0°C garage. What is the pressure of basketball outside where the temperature is 5.00°C? a. 0.375 atm b. 1.42 atm c. 1.54 atm d. 1.67 atm e. 1.58 atm 36. Consider the skeletal structure below. What is the chemical formula for this compound? a. CgHg02 b. CgH1002 OH c. CgHg02 d. C8H1002 e. C8H702
Answer #1

35. According to charle's law,

V1/T1 = V2/T2

T1 = 20 + 273.15 = 293.15°K

T2 = 5 + 273.15 = 278.15°K

Therefore, V1 = 1.05393*V2

According to Boyle's law,

P1*V1 = P2*V2

Therefore, 1.5*(1.05393*V2) = P2*V2

Cancelling out common terms,

P2 = 1.58 atm . . . . . . . . Ans. (Option E)

36. Option B. Structure is given below: